OpenStudy (anonymous):

ashton drops a golf ball. its height at any time is given by the equation f(x)=20-5x^2 where x is the time elasped in seconds. when will the ball hit the ground?

OpenStudy (anonymous):

set the height =0 and solve for x. So 0=-5x^2+20. -20=-5x^2 4=x^2 x=+-sqrt(2) since you only care about the positive root (no such thing as negative time, the answer is 2) Plug in to check -5(2^2)+20=0 -5(4)=-20 -20=-20

OpenStudy (anonymous):

thanks!